Kellen’s boat goes 12 mph. Find the rate of the current of the river if she can go 6 mi upstream in the same amount of time she can go 10 mi downstream.
Answer by scott8148(6628)

t = d / r ___ the times are equal

6 / (12 - c) = 10 / (12 + c)

"cross" multiplying ___ 72 + 6c = 120 - 10c